OpenStudy (anonymous):

Trig Identities Cos^2 x [(pi/2) - x] /cos x

OpenStudy (anonymous):

what would you like to know?

OpenStudy (anonymous):

I would like to know how to simplify the expression by using the fundamental identities

OpenStudy (loser66):

hey, is it \(\dfrac{cos^2(\pi/2 -x)}{cos x}\)??

OpenStudy (anonymous):

\[\cos ^{2}x[(\pi/2)-x] /cosx\]

OpenStudy (loser66):

so, why cos^2 x and cos x from numerator and denominator respectively cannot cancel out?

OpenStudy (loser66):

to get cos x (pi/2 -x)?

OpenStudy (anonymous):

cos(x - pi/2) = cos x * cos pi/2 + sin x * sin pi/2 cos(x - pi/2) = sin x

OpenStudy (anonymous):

cos(x - pi/2) = cos(pi/2 - x) because cosine is an even function. = sin(x) because co-functions of complements are equal.

OpenStudy (anonymous):

can cos^2x cross out with cos x even though they dont have the same exponent?

OpenStudy (anonymous):

1 sec let me think about that.

OpenStudy (anonymous):

i really don't know about that.

hartnn (hartnn):

cos^2x can cancel with cos x but do we have same angle for cos here ?? is your numerator cos x (pi/2 -x) or just cos (pi/2-x) ??

hartnn (hartnn):

\(\large \dfrac{\cos^2x}{\cos x} = \dfrac{\cos x \times \cancel {\cos x} }{\cancel{\cos x}}= \cos x\)

OpenStudy (anonymous):